OpenStudy (anonymous):

square root of (2x+3)-x= 0

OpenStudy (anonymous):

start with \[\sqrt{2x+3}=x\] the square both sides to get \[2x+3=x^2\] the \[x^2-2x-3=0\] \[(x-3)(x+1)=0\] so \[x=3\text{ or } x=-1\] then check both answers

OpenStudy (anonymous):

you will see that 3 works, but -1 does not